OpenStudy (anonymous):

Identify all values of t E (0,2pi] that satisfy tan(t)=-square root of 3 I need an example to help me solve this

OpenStudy (campbell_st):

well the angles you are looking for are in the 2nd quadrant \[\pi - \theta\] and 4th quadrant \[2\pi - \theta\] this is because these are the quadrants where tan is negative. you question involves and exact value  so find the answer for the postive have of tan from the triangle above then the answers are pi - answer 2pi - answer. hope it helps